OpenStudy (anonymous):

Solve the equations 2y=x+2 and 4x-8y=0

OpenStudy (anonymous):

r u sure this is correct qstn?

OpenStudy (anonymous):

bcz both eqns r of same kind.. x-2y=-2---(1) 4x-8y=0---(2) when eqn (1) is multiplied by 4 we got 4x-8y=-8 which is not correct

OpenStudy (luigi0210):

When you solve them what do you get?

OpenStudy (anonymous):

Hmm should this be set like -x + 2y - 2 =0?

OpenStudy (anonymous):

To where the answer should be y = x + 2/ 2

OpenStudy (anonymous):

Or writen as 

OpenStudy (anonymous):

oh ok thank you

OpenStudy (anonymous):

Your welcome :)